OpenStudy (anonymous):

which term in progression,18,-12,8,.....is 512/729

OpenStudy (anonymous):

Correction : Second Term is +12. ......................................… 18, 12, 8, 16/3, ... This is a G.P. with a = 18 r = 12/18 = 2/3. If the (n)th term is 512 / 729, then a. r^(n-1) = 512/729 18. r^(n-1) = 512/729 r^(n-1) = 512 / ( 729 x 18 ) = 256 / (27². 3²) = ( 2^8)/(3^8) = (2/3)^8 (2/3)^(n-1) = (2/3)^8 n - 1 = 8 n = 9 .................. Ans. Thus ( 9 )th term is 512/729.